# Tunneling in multi-site mesoscopic quantum Hall circuits

### v1

@@ VIII Heating effect in a multi-site QH circuit
Using the expression of the free correlator given in equation (5) followed by subtraction of corresponding vacuum contribution, the above equation provides the heat current carried by a free chiral edge at temperature $T$ as $J_{T}=\pi^{2}T^{2}/6R_{q}$. Thus, because of $J_{V}$ and $J_{T}$, in general, the grains get indeed heated. This heating is always opposed by heat dissipation to phonons in the sample. Due to the heating and subsequent colling effects on top of a complicated charging energy profile, the grains in general acquire a non-equilibrium temperature profile that can be predicted by solving the full kinetic equation [52, 53].
To get the rough estimates for the temperatures of the grains in the presence of applied voltage, we here follow a rather simplifying approach. We assume that in each grain, equilibrium is established at some unknown temperature $T_{j},j=1,2,3,4$, and neglected the back-action (cooling) produced by phonons. In this case, a set of four heat balance equations can be solved. The temperatures of the grains and that of the incoming edges $T$ in the setup 1 are accommodated in the correlators of the corresponding bosonic fields. Namely, the correlators $\left<\Phi_{j+1,\rm R}^{0}(\omega)\Phi_{j+1,\rm R}^{0}(\omega^{\prime})\right$> and $\left<\Phi_{j,\rm L}^{0}(\omega)\Phi_{j,\rm L}^{0}(\omega^{\prime})\right$> satisfy equation (5) with $T=T_{j}$, and similarly that for the incoming fields $\left<\Phi_{1,\rm R}^{0}(\omega)\Phi_{1,\rm R}^{0}(\omega^{\prime})\right$> and $\left<\Phi_{5,\rm L}^{0}(\omega)\Phi_{5,\rm L}^{0}(\omega^{\prime})\right$> at temperature $T$.
@@
To solve the heat balance equation, we also need to know the voltage distribution among the edge channels after biasing the field $\Phi_{\rm 1,R}$ by the voltage $V$. Such distributions can easily be obtained just by demanding that the current carried by the fields $\Phi_{\rm\alpha,R}-\Phi_{\rm\alpha,L}$ is the same for all $\alpha$. In this way, we obtain the voltage carried by the right and left moving edges as
@@
In the case where the two incoming edge channels described by the fields $\Phi_{\rm 1,R}$ and $\Phi_{\rm 5,L}$ are kept at different temperature $T_{s/d}$, the temperature in the above equation will be modified as $T^{2}\to(T^{2}_{s}+T^{2}_{d})/2$. [The asymptotic limit $T^{2}_{j}=(T^{2}_{s}+T^{2}_{d})/2$ is an expected result since, at the fully ballistic limit, the multi-site QH circuit can be thought of as an equivalent circuit representing a big composite grain coupled to two reservoirs. The latter is true under the assumption that the edge states between grains are fully covered by the connecting grains and are responsible for only proving impedances to the incoming electrons.] Similarly, in the case that the incoming edges are biased by a different voltage, $V_{s/d}$, the voltage in the equation (89) changes to $V\to V_{s}-V_{d}$. As seen from equation (89), in the out-of-equilibrium setting of setup 3, the heating effects are indeed significant. In all the linear response calculations presented earlier, the inclusion of the heating effect will give vanishingly small corrections that can possibly be neglected.
@@
We note that in the above calculations, we neglected the charge-charge coupling between the neighboring grains. In the realistic situations, such simplifying assumption however fails, and indeed an appreciable inter-island Coulomb interaction, that could even be of the same order of magnitude as that of intra-island, always exists [41]. The inter-island Coulomb interactions can be modeled by the cross-capacitances $C_{c}$. These capacitances do not allow the DC current to follow, but mediate an exchange of energy or heat between the neighboring grains. Therefore, the inter-island Coulomb interactions strongly affect the different aspects of heat transport in multi-site QH circuits, and must be taken into account to have any quantitative predictions for corresponding heat transport properties. At the technical label, when the cross-capacitance $C_{c}$ and direct-capacitance $C$ are both present, the estimation of heating effects will not lead to a simple result (89), since both capacitances can not tend to zero simultaneously [54]. One can also imagine a theoretical scenario where the grains are really far apart such that inter-grain interactions can be neglected. In this case, one is required to consider the finite segment of the circuit between the neighboring grains. The latter problem would then be identical to that of a short quantum wire connected to the leads having finite charging energies, and is known to have different physics than the usual QH circuits.
IX Conclusion
@@
We explored the transport properties of multi-site QH circuits, with special focus on the prototypical four-site setup where higher-order backscattering processes are the relevant perturbations. We showed that the consideration of such a higher order process is crucial in order to correctly identify the quantum critical features of the corresponding model. The critical properties are shown to be directly imprinted in the charge current flowing through the circuit in response to applied voltage bias. We investigated tunneling phenomena in the multichannel, multi-site mesoscopic QH circuits by explicitly evaluating their transport properties and verified their robust non-Fermi-liquid behaviors. We proposed a coveted pathway, based on looping the QH edge channels, to realize exotic zero temperature critical phenomena in multi-site QH circuits. We showed that the proposed method can realize numerous critical exponents by adjusting the number of grains and that of the looped edge channels. The effects of Joules heating in the explored devices are also uncovered.

### v2

@@ X Heating effects in a multi-site QH circuit
Using the expression of the free correlator given in equation (5), followed by subtraction of corresponding vacuum contribution, the above equation provides the heat current carried by a free chiral edge at temperature $T$ as $J_{T}=\pi^{2}T^{2}/6R_{q}$. Thus, because of $J_{V}$ and $J_{T}$, in general, the grains get indeed heated. This heating is always opposed by heat dissipation to phonons in the sample. Due to the heating and subsequent cooling processes, together with the complex charging energy profile, the grains generally acquire a nonequilibrium temperature distribution, which can be determined by solving the full kinetic equation [55, 56].
To obtain rough estimates for the grain temperatures in the presence of applied voltage, we here follow a rather simplifying approach. We assume that equilibrium is established within each grain at some unknown temperature $T_{j},j=1,2,3,4$, and neglect the back-action (cooling) produced by phonons. In this case, a set of four heat balance equations can be solved. The grain temperatures and the temperature $T$ of the incoming edges in setup 1 enter the correlators of the corresponding bosonic fields. Namely, the correlators $\left<\Phi_{j+1,\rm R}^{0}(\omega)\Phi_{j+1,\rm R}^{0}(\omega^{\prime})\right$> and $\left<\Phi_{j,\rm L}^{0}(\omega)\Phi_{j,\rm L}^{0}(\omega^{\prime})\right$> satisfy equation (5) with $T=T_{j}$, and similarly that for the incoming fields $\left<\Phi_{1,\rm R}^{0}(\omega)\Phi_{1,\rm R}^{0}(\omega^{\prime})\right$> and $\left<\Phi_{5,\rm L}^{0}(\omega)\Phi_{5,\rm L}^{0}(\omega^{\prime})\right$> at temperature $T$.
@@
To solve the heat-balance equations, we also need to know the voltage distribution among the edge channels after biasing the field $\Phi_{\rm 1,R}$ by the voltage $V$. Such distributions can easily be obtained just by demanding that the current carried by the fields $\Phi_{\rm\alpha,R}-\Phi_{\rm\alpha,L}$ is the same for all $\alpha$. In this way, we obtain the voltage carried by the right and left moving edges as
@@
In the case where the two incoming edge channels described by the fields $\Phi_{\rm 1,R}$ and $\Phi_{\rm 5,L}$ are kept at different temperatures $T_{s/d}$, the temperature in the above equation will be modified as $T^{2}\to(T^{2}_{s}+T^{2}_{d})/2$. [The asymptotic limit $T^{2}_{j}=(T^{2}_{s}+T^{2}_{d})/2$ is an expected result, since in the fully ballistic limit, the multi-site QH circuit can be thought of as an equivalent circuit representing a big composite grain coupled to two reservoirs. The interpretation holds true under the assumption that the edge states between grains are fully covered by the connecting grains and are responsible only for proving impedances to the incoming electrons.] Similarly, in the case that the incoming edges are biased by a different voltage, $V_{s/d}$, the voltage in the equation (89) changes to $V\to V_{s}-V_{d}$. As seen from equation (89), in the out-of-equilibrium setting of setup 3, the heating effects are indeed significant. In all the linear response calculations presented earlier, the inclusion of heating effects leads to vanishingly small corrections that can thus be neglected.
@@
We note that in the above calculations, we neglected the charge-charge coupling between the neighboring grains. In the realistic situations, such simplifying assumption however fails, and indeed an appreciable inter-island Coulomb interactions, that could even be of the same order of magnitude as that of intra-island, always exist [44]. The inter-island Coulomb interactions can be modeled by the cross-capacitances $C_{c}$. These capacitances do not allow the DC charge transport, but mediate an exchange of energy or heat between the neighboring grains. Therefore, inter-island Coulomb interactions strongly affect the different aspects of heat transport in multi-site QH circuits, and must be taken into account to obtain quantitative predictions for corresponding heat-transport properties. At the technical label, when the cross-capacitance $C_{c}$ and direct-capacitance $C$ are both present, the estimation of heating effects will not lead to a simple result (89), since both capacitances can not tend to zero simultaneously [57]. One can also imagine a theoretical scenario where the grains are really far apart such that inter-grain interactions can be neglected. In this case, one is required to consider the finite segment of the circuit between the neighboring grains. The latter problem is then equivalent to that of a short quantum wire connected to leads with finite charging energies, and is known to have different physics than the usual QH circuits.
XI Conclusion
@@
We explored the transport properties of multi-site QH circuits, with particular focus on the prototypical four-site setup, where higher-order backscattering processes constitute the relevant perturbations. We showed that consideration of such a higher order processes is crucial for correctly identifying the quantum critical features of the corresponding model. These critical properties are shown to be directly imprinted in the charge current flowing through the circuit in response to an applied voltage bias. We also investigated tunneling phenomena in the multichannel, multi-site mesoscopic QH circuits by explicitly evaluating their transport properties and demonstrated robust non-Fermi liquid behavior. We further proposed a coveted pathway, based on looping QH edge channels, to realize exotic zero-temperature critical phenomena in multi-site QH circuits. This approach allows the realization of a wide range of critical exponents by adjusting the number of grains and that of the looped edge channels. Finally, we quantified the effects of Joule heating in the explored devices. Our work paves the way for simulating a broad range of strongly correlated phenomena in nanoengineered electronic circuit with well-defined quantum Hall channels.